A coin lying on a meterstick starts to slide down when one end of the stick is raised



36.4cm above the horizontal. Find



a) The limiting angle for friction



b) The coefficient of starting friction

1
Expert's answer
2022-04-26T14:56:19-0400

a)

"\theta=\arctan\frac lm=20°,"

b)

"\mu=\frac lm=0.364."